# SounDiT: Geo-Contextual Soundscape-to-Landscape Generation

### v1

@@ 3.3 SounDiT
To support the GeoS2L task, we propose SounDiT, a diffusion-based generative model that synthesizes geographically coherent landscape images conditioned on environmental soundscapes and optional scene prompts (Figure 2). SounDiT advances geo-contextual landscape generation through two key innovations: (1) the SounDiT Block, which integrates sound, scene, and visual cues into the denoising process; and (2) a geo-contextual Retrieval-Augmented Generation (RAG) to guide image generation with additional geographic contexts at inference time.
Multimodal Encoders. To incorporate acoustic, semantic and visual features into SounDiT, we employ three encoder modules: (1) a Soundscape Encoder $\mathcal{E}_{\mathrm{Sound}}$, (2) a Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ and (3) a Landscape Encoder–Decoder pair $\bigl{(}\mathcal{E}_{L},\mathcal{E}^{\prime}_{L}\bigr{)}$. The Soundscape Encoder $\mathcal{E}_{Sound}$ embeds soundscape $s$ into an audio embedding $e_{\mathrm{sound}}$. The Scene Encoder $\mathcal{E}_{\mathrm{Scene}}$ encodes the optional scene prompt $c$ into a text embedding $e_{\mathrm{scene}}$. The Landscape Encoder $\mathcal{E}_{L}$ compresses a landscape image $l_{i}\in\mathbb{R}^{H\times W\times 3}$ into a latent embedding $e_{\mathrm{landscape}}$, and the Landscape Decoder $\mathcal{E^{\prime}}_{L}$ reconstructs the latent embedding back into a landscape image $\hat{l_{i}}\in\mathbb{R}^{H\times W\times 3}$.
SounDiT Block. Despite the demonstrated effectiveness of DiT in image generation, they have not been integrated into audio2image tasks. To effectively model geographic contexts from both soundscape and scene inputs, we propose the SounDiT block, a geo-contextual DiT to generate landscape images that align with intended geographic settings. The SounDiT operates in three stages: First, the timestep embedding $e_{\mathrm{t}}$ is fused with landscape latent embedding $e_{\mathrm{landscape}}$, and fed into a multi-head self-attention layer to capture visual features. Second, a multi-head cross-attention conditioned on the scene embedding $e_{\mathrm{scene}}$ is applied to enrich these features with scene semantics, producing scene-landscape representation. Third, to further integrate multi-level soundscape features, we introduce Time-Aware Prototype Mixture-of-Experts (TA-PMoE) module, composed of $M$ parallel cross-attention layers as experts. Each cross-attention layer fuses the scene-landscape representation with the soundscape embedding $e_{\mathrm{sound}}$, resulting in $M$ soundscape-scene-landscape representations $\mathbf{P}=[\,\mathbf{p}_{1},\dots,\mathbf{p}_{M}\,]$. A router $\mathbf{w}$ then calculates a weighted combination of these representations $\mathbf{P}$ based on temperature-scaled Softmax:
[…]
Here, $\tau>0$ controls the sharpness of the distribution: smaller values of $\tau$ concentrate weight on the most relevant soundscape-scene-landscape representation, while larger values lead to a smoother allocation across representations. Finally, the combined soundscape-scene-landscape representation is passed through a pointwise feedforward network to produce the denoised output. By hierarchically fusing visual, scene, and soundscape representations, SoundDiT effectively generates geo-contextually coherent landscape images.
Geo-contextual RAG Inference. In practice, additional geo-contextual information, such as location, scene and spatial layout [47], may be collected alongside soundscapes and landscapes to guide image generation. Thus, we propose a flexible geo-contextual RAG pipeline to further enhance the geographic coherence of the generated landscape images. Our method consists of two stages: (1) Geo-contextual Knowledge Database (GeoKD), where we store soundscape embeddings along with associated scene categories and layout representations with geographic locations; and (2) Geo-contextual Initialization, where the most relevant scene and layout are retrieved based on soundscape similarity and geographic proximity. These retrieved cues are then used to initialize the diffusion generation to enhance geographic coherence.
4 Datasets: SoundingSVI and SonicUrban
Prior audio2image studies have primarily relied on general datasets such as VEGAS [85], VGGSound [9], and ACAV100M [40]. Despite their proven effectiveness, these datasets lack geographic contexts. For instance, geographers may be more interested in generating a wetland implied by a frog call, rather than an image of the frog itself. A car horn might suggest a congested urban street, rather than a static vehicle. Some datasets, such as Landscape [41] and IntoTheWild [42], have included paired soundscape and image data, but offer limited spatial coverage and scene diversity. To address this gap, we introduce two new multimodal datasets: SoundingSVI and SonicUrban, designed to support GeoS2L generation with large-scale, geographically diverse, and context-rich data. Figure 3 provides several sample images of these two datasets.
[…]
Dataset Modalities # Size Geo-Contextual Scene Diversity Spatial Coverage Viewpoint
ACAV100M [40] Video 100M -- -- -- Ground-level
VEGAS [85] Video 28K -- -- -- Ground-level
[…]
IntoTheWild [42] Video 142 $\checkmark$ Low -- Ground-level
Landscape [41] Video 1747 $\checkmark$ Medium -- Ground-level
SoundingEarth [28] Audio & Image 50K $\checkmark$ High High Overhead
SoundingSVI (Ours) Audio & Image 330K $\checkmark$ High High Ground-level
SonicUrban (Ours) Video 250K $\checkmark$ High High Ground-level
Figure 3: Representative samples from the SoundingSVI and SonicUrban datasets, showing diverse geographic contexts.
SoundingSVI. The SoundingSVI dataset was constructed from geotagged soundscape recordings from the Aporee platform [56]. For each raw soundscape recording, we retrieved nearby Google Street View Image (SVI) using its latitude and longitude to capture the surrounding visual landscapes. This follows a similar approach to the Sounding Earth dataset [28], which paired geotagged audio with remote sensing imagery. The raw recordings were then segmented into 10-second clips, with a human voice detection model [73] used to filter clips dominated by human voice. Notably, multiple candidate SVIs captured from different angles and timestamps might be retrieved for clips from the same recording. A sound source localization model [58] was further applied to match each clip with the most relevant landscape image. In total, SoundingSVI dataset comprises 334,495 soundscape-landscape pairs in 90 countries. To support optional scene prompts during training, a scene type for each soundscape-landscape pair was predicted by a scene classification model [27].
SonicUrban. The SonicUrban dataset leverages unedited videos as a source of temporally and spatially consistent soundscape-landscape pairs. Following a prior practice [86], we searched and manually checked YouTube videos that contain real-world soundscapes and landscapes, using a combination of city names and keywords, like “City walk; New York”. Each video was processed into 10-second audio clips, and 10 evenly distributed frames were extracted for each clip. The same human voice detection model [73] and sound source localization model [58] used in SoundingSVI were applied to filter human-voice dominated clips and identify the most representative frame for each audio clip, respectively. In total, the constructed SonicUrban consists of 250,310 soundscape-landscape pairs across 131 cities and 97 countries. A scene classification model [27] was further employed to generate a scene prompt for each pair as a geographic scene condition.
5 Geo-Contextual Evaluation Framework
Figure 4: Geo-contextual evaluation framework. The ground truth and output landscape images are assessed at the element-, scene-, and human-perception levels following geography and urban planning practices.
Prior audio2image studies commonly rely on evaluation metrics such as Fréchet Inception Distance (FID) [30], Audio-Image Similarity (AIS) [30], and Image-Image Similarity (IIS) [5], to assess visual or audio fidelity, overlooking the underlying geographic settings and target scenes. Since soundscapes and landscapes co-occur in the same space, they are expected to share similar environmental characteristics and place settings. Following this hypothesis, we introduce a novel practically-informed geo-contextual evaluation framework — Place Similarity Score (PSS) — that measures the environmental settings of soundscape and landscape across element, scene, and human perception levels (Figure 4). These three levels of measurements of environmental characteristics have been widely employed in geographic and urban planning practices [50, 36].
Table 2: Place Similarity Score (PSS)
Score Definition
$\text{PSS}_{\text{element}}$ $\frac{1}{n}\sum_{i=1}^{n}\frac{E_{i}\cdot\hat{E}_{i}}{\|E_{i}\|\|\hat{E}_{i}\|}$
$\text{PSS}_{\text{scene}}$ $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{P_{i}^{\text{k}}\cap T_{i}^{\text{k}}\neq\varnothing\right\}$
$\text{PSS}_{\text{perception}}$ $\frac{1}{n}\sum_{i=1}^{n}\left\|R(l_{i})-R(\hat{l}_{i})\right\|_{1}$
Element-level Evaluation quantifies the similarity of geographic elements (e.g., trees [43], sky [23], water bodies [69], traffic signs [3], and buildings [13]) between the geographical contexts of the input soundscape and generated landscapes [35]. We employ the DeepLabV3 semantic segmentation model [12], pre-trained on the ADE20K dataset [84], to identify 150 predefined geographic elements $E_{i}$ in each image $i$. For each image $i$, we calculate a normalized element distribution vector $E_{i}=\{e_{i}^{1},e_{i}^{2},\dots,e_{i}^{k}\}$, where each $e_{i}^{k}$ represents the ratio of pixels identified as the $k$-th element. [17]. We calculate element ratio vectors $E_{i}$ and $\hat{E_{i}}$ from the ground truth landscape image $l_{i}$ and their corresponding image $\hat{l_{i}}$. We then measure the cosine similarity between each pair of real and generated images and report the average across all samples $n$ as the $PSS_{element}$.
Scene-level Evaluation is designed to evaluate the consistency of the entire environmental scene (e.g., “forest,” “beach,” or “residential area”) between the generated landscape images and their corresponding ground truth images [35]. To infer scene categories, we employ a ResNet50 model [27], pre-trained on the Places365 dataset [83], to identify 365 predefined scene categories. For each soundscape $s$, we predict the top-$k$ scene labels and evaluate whether the generated image $\hat{l_{i}}$ aligns with the same scene category as its ground truth image $l_{i}$. The $PSS_{scene}$ is then calculated by its top-$k$ accuracy, where $k=1$ or $5$. $n$ is the total number of images, $P_{i}^{\text{k}}$ denotes the top-$k$ predicted category set for the $i$-th generated image, $T_{i}^{\text{k}}$ denotes the top-$k$ ground truth category set for the $i$-th real image.
Human Perception-level Evaluation is designed to evaluate whether the generated and real landscape images evoke similar human subjective perceptions and feelings, such as whether a place makes people feel “safe” or “lively” [82]. We utilize a DenseNet121 model [31], pre-trained on the MIT Place Pulse dataset [55], to measure six dimensions of human perceptions of environment, including “Safe,” “Beautiful,” “Depressing,” “Lively,” “Wealthy,” and “Boring”. For each image $i$, we compute the perception scores in six dimensions for the generated image $\mathrm{R}(\hat{l}_{i})$ and its corresponding ground truth $\mathrm{R}(l_{i})$. $PSS_{perception}$ is then defined as the average L1 distance across six dimensions, where $n$ is the number of the test set, $R$ is the Place Perception scores of each dimension.
By integrating metrics across element-, scene-, and human perception levels (Table 2), we provide a holistic geo-contextual evaluation framework. Informed by geographic domain knowledge and urban planning practices, we extend beyond measuring visual quality to assessing whether generated landscape images are contextually aligned with the environmental characteristics in input soundscapes.
6 Experiments
Figure 5: Visual comparison of landscape images generated by baseline models (CoDi, Sound2Scene, AudioToken Pre-trained/SD1/SD2, GlueGen) and our proposed SounDiT across SoundingSVI, SonicUrban, and two benchmark datasets. Ground truth landscape images are provided in the last column. Our proposed SounDiT captures significantly more geographic contexts than baselines.
6.1 Implementation Details
Model Architecture and Training. We adopt the pre-trained Variational Autoencoder (VAE) modules from Latent Stable Diffusion [65] as our landscape encoder–decoder pair, which were originally trained on the COCO dataset [45]. For soundscape conditioning, we employ the ImageBind-Huge audio encoder [22], pre-trained on two million AudioSet clips [21] covering a wide range of environmental soundscapes. We integrate a scene encoder based on the T5 architecture [62], which transforms high-level semantic scene prompts into latent embeddings to provide geographic semantics. Our SounDiT Block is initialized with the weights of PixArt-$\alpha$, a pre-trained text-to-image DiT model [10], and four experts are employed in TA-MoE Module. All components are trained using a learning rate of $1\times 10^{-4}$. Experiments are conducted on NVIDIA A100 and A6000 GPUs.
Benchmark Datasets. We curated two additional GeoS2L benchmark datasets to evaluate model generalizability. The first is a geo-contextual subset of VGGSound [9], comprising 16 classes with environmental semantics. It contains 6,713 soundscape–landscape pairs. The second combines the Landscape [41] and IntoTheWild [42] datasets, resulting in 1,889 soundscape-landscape pairs.
Baseline Models. We selected four state-of-the-art audio2image models, including CoDi [72], Sound2Scene [70], AudioToken [80], and GlueGen [61] as baseline models. For CoDi, we utilized its official pre-trained model. For AudioToken, we employed the pre-trained model without additional training, and trained it on our datasets using Stable Diffusion 1 and 2 as its backbone, respectively. The other two baselines are training from scratch based on their released code and training parameters.
Benchmark Evaluation Metrics. To evaluate the model performance, we adopted FID [30], AIS [30], and IIS [5], effective in previous audio-conditioned image generation studies [5, 71]. FID compares Gaussian distributions of features extracted from real and generated images, using a pre-trained Inception model. AIS measures audio-image semantic similarity using Wav2CLIP [78]. IIS evaluates image-image semantic similarity using CLIP-based embedding [29].
6.2 Main Results
[…]
General Metrics Place Similarity Score
@@ 6.2 Main Results
Sound2Scene 78.137 0.586 0.623 0.467 0.374 0.735

### v2

@@ 3.3 SounDiT
To effectively and efficiently inject soundscape and scene context, each SounDiT block introduces (1) a Mixture-of-Experts (MoE) Soundscape Conditioning module, (2) a Scene Low-Rank Mixer (SLRCM) module, and a Scene AdaLN (S-AdaLN) module to improve scene consistency.
Multimodal Encoders. Following latent diffusion models (LDM) [66], landscape images are reconstructed within the latent space. Given a landscape image $l\!\in\!\mathbb{R}^{H\times W\times 3}$, a VAE encoder–decoder $(\mathcal{E}_{L},\mathcal{D}_{L})$ is employed to encode it into the latent embedding $\mathbf{e}_{l}=\mathcal{E}_{L}(l)\in\mathbb{R}^{h\times w\times c}$ and reconstruct the latent embedding back into a image $\hat{l}=\mathcal{D}_{L}(\mathbf{e}_{l}).$ To integrate acoustic features and geo-contextual semantics into denoising process, a multi-modal encoder is adopted to encode the soundscapes $s$ and the scene prompt $c$ into a shared latent space, resulting in a soundscape embedding $\mathbf{e}_{s}$ and a scene embedding $\mathbf{e}_{c}$.
SounDiT block. The SounDiT block operates in four stages. First, we apply multi-head self-attention layer conditioned by the timestep embedding $e_{\mathrm{t}}$ via AdaLN-Zero [26], preserving compatibility with pretrained DiT backbones. Second, we design a lightweight Scene Low Rank Context Mixer (SLRCM) module to inject scene context. SLRCM introduces a low-rank residual path inside each SounDiT block. Given an input token $x\in\mathbb{R}^{B\times N\times D}$ and scene embedding $\mathbf{e}_{c}\in\mathbb{R}^{B\times D_{c}}$, SLRCM modulates tokens through a rank-$r$ linear operator parameterized by $\mathbf{e}_{c}$. We specify the low-rank content mixer $A(\mathbf{e}_{c})$ that transforms token features along an $r$-dimensional path, together with a sample-wise strength $s(\mathbf{e}_{c})$ that controls the overall contribution of this path:
$A(\mathbf{e}_{c})=W_{q}\,\mathrm{Diag}\!\big(\tanh(\phi(\mathbf{e}_{c}))\big)\,W_{v},$ (5)
where $W_{q}\!\in\!\mathbb{R}^{D\times r}$ and $W_{v}\!\in\!\mathbb{R}^{r\times D}$ are learned low-rank projections. The mapping $\phi:\mathbb{R}^{D_{c}}\!\to\!\mathbb{R}^{r}$ converts the scene embedding into an $r$-dimensional gating vector, and the diagonal operator applies this gate element-wise along the rank-$r$ channel, yielding a compact, diagonally gated low-rank transform. In addition, we define a positive, sample-wise strength to regulate the residual magnitude:
$s(\mathbf{e}_{c})=g(\mathbf{e}_{c})\,\mu\,\tanh(\alpha),$ (6)
where $g:\mathbb{R}^{D_{c}}\!\to\!\mathbb{R}_{>0}$ produces a per-sample scale (e.g., via softplus), $\mu$ is a global guidance scalar, and $\alpha$ is a learnable scalar bounded. Given these definitions, the token update becomes:
$x^{\prime}\;=\;x\;+\;s(\mathbf{e}_{c})\,\mathrm{LN}(x)\,A(\mathbf{e}_{c}),$ (7)
where $\mathrm{LN}(\cdot)$ denotes layer normalization. SLRCM thus provides a compact, diagonally gated low-rank modulation that preserves the pre-trained attention structure while injecting scene priors with low computational costs.
Third, to further integrate multi-level soundscape features, we introduce MoE Soundscape Conditioning module, composed of $M$ experts that share the key $K_{s}$ and value $V_{s}$ while maintaining expert-specific, low-rank query $Q$. Given input token $x\in\mathbb{R}^{B\times N\times D}$,we compute $K_{s}=s\,W_{K},V_{s}=s\,W_{V}$ once and for each expert $m\in\{1,\dots,M\}$ form a low-rank query $Q_{m}$
$Q_{m}=f(x)\,W^{(m)}_{q\downarrow}W^{(m)}_{q\uparrow}\!$, (8)
where $f$ is a per-token Layer Norm. Each expert then produces
$Z_{m}\;=\;\mathrm{MHA}\bigl(Q_{m},\,K_{s},\,V_{s}\bigr)\,W_{O},$ (9)
where $W_{O}$ is shared across experts, $Z_{m}$ refers to soundscape-conditioned token features, and $MHA$ is multi-head attention.
Routing weights are computed by a temperature-scaled dot product between the audio summary and learnable prototypes $\mathbf{P}\in\mathbb{R}^{d\times M}$ with time augmentation introduced through a projected timestep embedding:
$w=\operatorname{Softmax}\!\Bigl(\tfrac{1}{\tau}\;\mathbf{e}_{s}^{\top}\bigl(\mathbf{P}+\mathbf{W}_{t}\,e_{t}\,\mathbf{1}^{\top}\bigr)\Bigr),$ (10)
where $e_{t}$ is the timestep embedding projected by $\mathbf{W}_{t}$. Finally, we aggregate the experts via a top-$k$ soft mixture and apply a global audio gate to update tokens in each SounDiT block:
$x^{\prime}\;=\;x\;+\;\gamma\sum_{m\in\mathcal{K}}\operatorname{softmax}\!\bigl(w_{m}/\tau_{m}\bigr)_{m}\;Z_{m},$ (11)
where $\gamma=\tanh(\mathbf{e}_{s})$ is a scalar gate, yielding a smooth, zero-centered bounded activation. By enabling expert specialization under a fixed K/V budget, the MoE provides scalable conditioning that improves geo-contextual consistency across diverse soundscapes.
Finally, to enhance geo-contextual consistency, we introduce a practical-informed S-AdaLN, an AdaLN-Zero extension, which derives scale–shift parameters from the timestep embedding $\mathbf{e}_{t}$ and the scene embedding $\mathbf{e}_{c}$ via learnable bounded mixing. The modulated tokens are processed by a pointwise feed-forward network, followed by a gated residual add, yielding the block output for noise-residual prediction. By hierarchically fusing visual, scene, and soundscape cues, SounDiT generates geo-contextually coherent landscapes.
4 Datasets
Table 1: Comparison of audio-visual datasets. $\dagger$} denotes partial/limited geo-context.
Dataset Modalities # Size Geo-Contextual Viewpoint
ACAV100M [44] Video 100M - Ground
VEGAS [87] Video 28K - Ground
VGGSound [10] Video 200K $\dagger$} Ground
IntoTheWild [46] Video 142 Ground
Landscape [45] Video 1747 Ground
SoundingEarth [30] Audio & Image 50K Overhead
@@ 4 Datasets
Prior A2I studies have primarily relied on large-scale, general-domain audio-visual datasets, such as VEGAS [87], VGGSound [10], and ACAV100M [44]. These datasets primarily focus on sound-source localization and general audio-visual correspondence learning, lacking the specific geographical context required for soundscape-to-landscape generation. For instance, geographers may be more interested in generating a wetland implied by a frog call, rather than an image of the frog itself. A car horn might suggest a congested urban street, rather than a static vehicle. Some datasets, such as Landscape [45] and IntoTheWild [46], have included paired soundscape and image data, but offer limited spatial coverage and scene diversity. To address this gap, we introduce two new multimodal datasets: SoundingSVI and SonicUrban, designed to support GeoS2L generation with large-scale, geographically diverse, and context-rich data.
SoundingSVI. The SoundingSVI dataset was constructed from geotagged soundscape recordings from the Aporee platform [59]. For each raw soundscape recording, we retrieved nearby Google Street View Image (SVI) using its latitude and longitude to capture the surrounding visual landscapes. This follows a similar approach to the Sounding Earth dataset [30], which paired geotagged audio with remote sensing imagery. The raw recordings were then segmented into 10-second clips, with a human voice detection model [74] used to filter clips dominated by human voice. Notably, multiple candidate SVIs captured from different angles and timestamps might be retrieved for clips from the same recording. A sound source localization model [61] was further applied to match each clip with the most relevant landscape image. To ensure temporal consistency and account for environmental changes, we refined the dataset by discarding soundscape-SVI pairs with significant time gaps. Furthermore, to obtain annotated place semantics, a Vision-Language Model (VLM), Qwen2.5-VL-7b, [2] was used to annotate each pair with a scene context (e.g., street and residential neighborhood), which serve as the scene prompts for our model. In total, SoundingSVI dataset comprises 169,221 soundscape-landscape pairs in 90 countries.
SonicUrban. The SonicUrban dataset leverages unedited videos as a source of temporally and spatially consistent soundscape-landscape pairs. Following a prior practice [89], we searched and manually checked YouTube videos that contain real-world soundscapes and landscapes, using a combination of city names and keywords, like “City walk; New York”. Each video was processed into 10-second audio clips, and 10 evenly distributed frames were extracted for each clip. The same human voice detection model [74] and sound source localization model [61] used in SoundingSVI were applied to filter human-voice dominated clips and identify the most representative frame for each audio clip, respectively. In total, the constructed SonicUrban consists of 236,674 soundscape-landscape pairs across 131 cities and 97 countries. The same VLM [2] was further employed to generate a scene prompt for each pair as a geographic scene condition.
5 Geo-Contextual Evaluation Suite
Prior A2I studies commonly rely on evaluation metrics such as Fréchet Inception Distance (FID) [33], Audio-Image Similarity (AIS) [33], and Image-Image Similarity (IIS) [5], to assess visual or audio fidelity. However, such metrics do not assess whether the generated image is coherent with the underlying geographic setting or the intended target scene. Since soundscapes and landscapes co-occur in the same space, they are expected to share similar environmental characteristics and place settings. Following this hypothesis, we introduce Place Similarity Score (PSS), a practically informed geo-contextual evaluation that quantifies the underlying place setting reflected in images across the element-, scene-, and human-perception levels. These three levels of measurements of environmental characteristics have been widely employed in geographic and urban planning practices [53, 41].
Element-level PSS. This metric measures the agreement in geographic elements (e.g., trees [47], sky [24], water bodies [70], traffic signs [3], and buildings [14]) between the ground-truth and generated landscapes [40]. We use DeepLabV3 [13] pre-trained on ADE20K [86] to segment K=150K\!=\!150 predefined elements. For each image $i$, let $\mathbf{e}_{i}\in\mathbb{R}^{K}$ and $\hat{\mathbf{e}}_{i}\in\mathbb{R}^{K}$ be the normalized element-ratio vectors derived from the ground-truth image $l_{i}$ and the generated image $\hat{l}_{i}$, respectively. Formally, The element-level score is calculated as:
$\mathrm{PSS}_{\text{elem}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\mathbf{e}_{i}^{\top}\,\hat{\mathbf{e}}_{i}}{\|\mathbf{e}_{i}\|_{2}\,\|\hat{\mathbf{e}}_{i}\|_{2}},$ (12)
where $\mathbf{e}_{i}=(e_{i1},\dots,e_{iK})^{\top}$ and $\hat{\mathbf{e}}_{i}=(\hat{e}_{i1},\dots,\hat{e}_{iK})^{\top}$; $n$ is the number of evaluation images.
Scene-level PSS. This metric is designed to evaluate the consistency of the entire environmental scene (e.g., forest, beach, or residential area) between the generated landscape images and their corresponding ground truth images [40]. To infer scene categories, we employ a ResNet50 model [29], pre-trained on the Places365 dataset [85], to identify 365 predefined scene categories. For each soundscape $s$, we predict the top-$k$ scene labels and evaluate whether the generated image $\hat{l_{i}}$ aligns with the same scene category as its ground truth image $l_{i}$.
$\mathrm{PPS}_{\text{Scene}}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{P_{i}^{\text{k}}\cap T_{i}^{\text{k}}\neq\varnothing\right\},$ (13)
where $k=1$ or $5$. $n$ is the total number of images, $P_{i}^{\text{k}}$ denotes the top-$k$ predicted category set for the $i$-th generated image, $T_{i}^{\text{k}}$ denotes the top-$k$ ground truth category set for the $i$-th real image.
Human Perception-level PSS.This metric is designed to evaluate whether the generated and real landscape images evoke similar human subjective perceptions and feelings, such as whether a place makes people feel “safe” or “lively” [83]. We utilize a DenseNet121 model [34], pre-trained on the MIT Place Pulse dataset [58], to measure six dimensions of human perceptions of environment, including safe, beautiful, depressing, lively, wealthy, and boring. For each image $i$, we compute six-dimensional place-perception scores for the ground-truth image $l_{i}$ and the generated image $\hat{l}_{i}$. The perception-level PSS is defined as:
$\mathrm{PSS}_{\text{perc}}=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathbf{R}(l_{i})-\mathbf{R}(\hat{l}_{i})\right\|_{1},$ (14)
where $\mathbf{R}(\cdot)\in\mathbb{R}^{6}$ denotes the six-dimensional perception score vector, and $n$ is the number of test images.
By integrating metrics across element-, scene-, and human perception levels, we provide a holistic geo-contextual evaluation framework. Informed by geographic domain knowledge and urban planning practices, we extend beyond measuring visual quality to assessing whether generated landscape images are contextually aligned with the environmental characteristics in input soundscapes.
Table 2: Quantitative evaluation results across SounDiT and baseline models on two proposed datasets and benchmark datasets. $\dagger$} indicates pre-trained models. We compare our SounDiT model with baselines using general metrics (FID, AIS, IIS) and our PSS metrics.
General Metrics Place Similarity Score (PSS)